40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jul 19 18:41:42 2022, yuta, Update, BHD, Contrast measurements for Michelson and ITM-LO ContrastMeasurements.pdf
    Reply  Wed Jul 20 18:07:52 2022, Paco, Update, BHD, BHD MICH test BHD_MICH_OSC.pdf
       Reply  Tue Jul 26 20:54:08 2022, Paco, Update, BHD, BHD MICH test - LO phase control 20220726_BICHD.pdf
          Reply  Wed Jul 27 18:30:50 2022, yuta, Update, BHD, LO beam power at BHD DCPDs is significantly lower than expected LOBeamAtBHD.JPGITMXSingleBounceAtBHD.jpg
             Reply  Thu Jul 28 14:34:40 2022, Yehonathan, Update, BHD, LO beam power at BHD DCPDs is significantly lower than expected 
                Reply  Thu Jul 28 16:51:55 2022, Tega, Update, BHD, Shaking test for LO beam AS beam to BHD DCPDs Screenshot_2022-07-28_16-58-34_LOandASShaking.pngScreenshot_2022-07-28_18-08-55_DCPDPOPSuspensionCoherence.png
                   Reply  Thu Jul 28 20:16:26 2022, Anchal, Update, BHD, Shaking test for LO beam AS beam to BHD DCPDs 
                   Reply  Fri Jul 29 18:24:53 2022, Tega, Update, BHD, LO beam power improved by factor of 6 after LO and AS beam alignment LO_and_oldAS_settings.pngBHD_fringe_settings.pngScreenshot_2022-07-29_19-37-39.pngFromTheLeft-AS-POP-LO.JPG
                      Reply  Wed Aug 3 16:00:51 2022, yuta, Update, BHD, BHD fringe aligned with reduced LO and AS beam clipping Screenshot_2022-08-03_15-46-36_BHDfringeAlmostUnclipped.png
                         Reply  Tue Aug 9 15:33:12 2022, yuta, Update, BHD, BHD fringe contrast improved from 43% to 74% HPC-DCPD_B_OUT_1344118517_ITMY-LO.pngHPC-DCPD_A_OUT_1344118517_ITMY-LO.pngHPC-DCPD_B_OUT_1344118318_ITMX-LO.pngHPC-DCPD_A_OUT_1344118318_ITMX-LO.png
                            Reply  Tue Aug 9 15:50:22 2022, Koji, Update, BHD, BHD fringe contrast improved from 43% to 74% 
                               Reply  Wed Nov 16 15:09:08 2022, yuta, Update, BHD, BHD fringe contrast measured with unwhitening filters ContrastMeasurements20221116_edited.pdf
                         Reply  Tue Oct 25 08:25:00 2022, JC, Update, BHD, BHD fringe aligned with reduced LO and AS beam clipping 
                         Reply  Thu Nov 17 11:24:39 2022, JC, HowTo, LSC, Locking MICH 
                            Reply  Tue Nov 29 11:38:37 2022, JC, HowTo, LSC, Lock Single Arm After MICH lock Screen_Shot_2022-11-29_at_11.01.56_AM.png
Message ID: 17067     Entry time: Tue Aug 9 15:33:12 2022     In reply to: 17056     Reply to this: 17068
Author: yuta 
Type: Update 
Category: BHD 
Subject: BHD fringe contrast improved from 43% to 74% 

[Anchal, Yehonathan, Yuta]

We did the constrast measurement with the method same as 40m/17020.
Contrast between ITM single bounce and LO beam increased to 74% (we had 43% before unclipping LO beam in 40m/17056).
From equations in 40m/17041 and measured ITM sigle bounce power (93 or 138 counts @ BHD DCPD) and LO power (130 or 124 counts @ BHD DCPD) from 40m/17056,  expected visibility for perfectly mode-matched case is 99%.
Measured constrast of 74% indicate mode-matching of 56%.

Both arms locked, MICH fringe (20% percentile)
Contrast measured by C1:LSC-ASDC_OUT is 80.66 +/- 0.20 %
Contrast measured by C1:LSC-POPDC_OUT is 92.27 +/- 0.66 %
Contrast measured by C1:LSC-REFLDC_OUT is 89.59 +/- 0.84 %
Contrast measured by all is 87.51 +/- 1.69 %

Both arms misaligned, MICH fringe (20% percentile)
Contrast measured by C1:LSC-ASDC_OUT is 82.50 +/- 0.61 %
Contrast measured by C1:LSC-POPDC_OUT is 94.18 +/- 0.26 %
Contrast measured by C1:LSC-REFLDC_OUT is 92.78 +/- 0.19 %
Contrast measured by all is 89.82 +/- 1.75 %

ITMX-LO fringe (40% percentile)
Contrast measured by C1:HPC-DCPD_A_OUT is 73.93 +/- 1.52 %
Contrast measured by C1:HPC-DCPD_B_OUT is 73.56 +/- 1.22 %
Contrast measured by all is 73.74 +/- 0.98 %

ITMY-LO fringe (40% percentile)
Contrast measured by C1:HPC-DCPD_A_OUT is 73.45 +/- 0.61 %
Contrast measured by C1:HPC-DCPD_B_OUT is 75.27 +/- 0.50 %
Contrast measured by all is 74.36 +/- 0.54 %

Attachment 1: HPC-DCPD_B_OUT_1344118517_ITMY-LO.png  17 kB  Uploaded Tue Aug 9 16:35:25 2022  | Hide | Hide all
HPC-DCPD_B_OUT_1344118517_ITMY-LO.png
Attachment 2: HPC-DCPD_A_OUT_1344118517_ITMY-LO.png  17 kB  Uploaded Tue Aug 9 16:35:34 2022  | Hide | Hide all
HPC-DCPD_A_OUT_1344118517_ITMY-LO.png
Attachment 3: HPC-DCPD_B_OUT_1344118318_ITMX-LO.png  17 kB  Uploaded Tue Aug 9 16:35:49 2022  | Hide | Hide all
HPC-DCPD_B_OUT_1344118318_ITMX-LO.png
Attachment 4: HPC-DCPD_A_OUT_1344118318_ITMX-LO.png  17 kB  Uploaded Tue Aug 9 16:36:06 2022  | Hide | Hide all
HPC-DCPD_A_OUT_1344118318_ITMX-LO.png
ELOG V3.1.3-